# Multi-Step Knowledge Interaction Analysis via Rank-2 Subspace Disentanglement

### v1

@@ 4.1 Experimental Setting
Dataset and Model: We conduct experiments on three open-weight decoder-only instruct-based LMs: Llama-3.1-8B (Meta-Team, 2024), Gemma-2 9B (Gemma-Team, 2024) and Mistral-v0.3 7B (Jiang et al., 2023) using four publicly available QA datasets: BaseFakepedia, MultihopFakepedia Minder et al. (2025), StrategyQA Geva et al. (2021), and OpenBookQA Mihaylov et al. (2018); Cheng et al. (2024). BaseFakepedia is a knowledge conflict dataset containing queries with 23 relations from Wikipedia. MultihopFakepedia is an extension of BaseFakepedia containing queries that require extra-hop reasoning in generating answers. StrategyQA is a multi-hop reasoning QA dataset with implicit reasoning steps present in the queries. OpenBookQA is a commonsense reasoning-based QA dataset.
@@
Explained Variance (Cumulative): Consider the matrix $H=[\{{\vec{\mathbf{h_{a_{j}}}}}\}_{j=1}^{N}],H\in\mathbb{R}^{d*d}$ as the concatenation of hidden representation $\vec{\mathbf{h_{a}}}\in\mathbb{R}^{d}$ of the answer token $a$ over $N$ examples. The $d$ singular values $\sigma_{1}>>\sigma_{2},...>>\sigma_{d}$ from the diagonal matrix $\Sigma\in\mathbb{R}^{d*d}$ after the Singular Value Decomposition (SVD) of $HH^{T}=A\Sigma B$ indicates the strength of $HH^{T}$ in $d$ orthonormal directions. For top-$r$ singular values $\sigma_{1}>>\sigma_{2},...>>\sigma_{r}$, the cumulative explained variance $EV_{r}$:
$EV_{r}=\nicefrac{{\sum_{j=1}^{r}\sigma_{j}^{2}}}{{\sum_{j=1}^{d}\sigma_{j}^{2}}}$ (8)
@@
Figure 2: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four QA datasets using Mistral-7B-Instruct-v0.3. The split noise denotes cases where answers from individual knowledge sources agree with each other but differ from the final answer, i.e., $a(q,c)=a(a,\varepsilon)$ and $a(a,c)\neq a$.
@@
Fig. 2 illustrates the Kernel Density Estimate plot of subspace components from the Mistral-7B-Instruct-v0.3 model, indicating the histogram distribution of different types of knowledge interactions smoothed using a Gaussian kernel parameterised by the kernel width as a hyperparameter.
Across all datasets and knowledge interaction types (§3.2), the mean of the subspace component distribution converges to 0, indicating significant contribution from the orthogonal complement $\mathbf{(I-P)}$, showing insufficiency of the rank-1 projection subspace $\mathbf{P}$ in encoding the PK-CK interaction. This observed behaviour for the ‘complementary’ and ‘supportive’ interactions supports our hypothesis that, for examples where both PK and CK contribute equally, the rank-1 $\mathbf{P}$ fails to distinctly encode individual contributions. However, surprisingly, we observe a similar behaviour for the ‘conflicting’ interaction type as well, illustrating that $\mathbf{P}$ fails to also differentiate where the conflicts arise from. We observe similar results for the other two models (see App. A.2). These results indicate that different knowledge interactions are poorly captured by the rank-1 projection subspace, with most of the interaction signal residing in the orthogonal complement $\mathbf{(I-P)}$, thereby suggesting the necessity of a higher-rank subspace to effectively disentangle PK–CK contributions.
To understand the minimum rank required for the learnt projection subspace where the orthogonal complement contribution is minimum and it encodes different knowledge interaction types, we plot the cumulative explained variance $EV_{r}$ for different ranks across all datasets and models in Fig. 3. $EV_{r}$ reaches $1.0$ at rank-2 and converges thereafter for all datasets and models.
@@
Figure 4: Patchscope on OpenBookQA from Meta-Llama-3.1-8B-Instruct. a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$. b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$.
Figure 5: Individual PK-CK contribution in generating the answer token $a$ for all the datsets from Meta-Llama-3.1-8B-Instruct model.
Figure 6: Distribution of different knowledge interaction types present in the dataset.
(a) Overall
(b) Supportive
(c) Complementary
(d) Conflicting
In this section, we investigate the dynamics of individual PK and CK contribution ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ at every sequence step $i$ over the NLE $\mathcal{E}$ generation. Among the four datasets, OpenBookQA shows the largest jump in cumulative explained variance from rank-1 to rank-2 across all models (Fig. 3). We therefore use OpenBookQA as our development set to (i) locate intent-sensitive layers via activation patching and (ii) learn the rank-2 projection subspace. We then freeze these choices and evaluate the resulting rank-2 subspace on all datasets. We begin by identifying important layers for encoding individual knowledge contributions towards the final answer generation in the rank-2 projection subspace. Fig. 4 shows that patching from BOTH→PRI (L13–18) yields a larger probability gain than BOTH→CTX (L15–17), indicating that Llama-3.1-8B-Instruct relies more on PK than on CK for OpenBookQA. We observe similar results for gemma-2-9b-it and Mistral-7B-Instruct-v0.3 (see Fig. 15, 16, App). This PK dominance in a context-rich task suggests that the model tends to recall commonsense information from memory rather than grounding answers in the provided context, highlighting potential for future work on context-sensitive knowledge control.
@@
To identify the individual knowledge contribution in generating the final answer token $a$, we project its hidden representation $\vec{\mathbf{h_{a}}}$ to the rank-2 projection subspace to compute the ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ as the PK-CK contributions. Fig. 5 indicates an overall higher CK contribution for the BaseFakepedia and MultihopFakepedia, and higher PK contribution for the StrategyQA and OpenBookQA for the Meta-Llama-3.1-8B-Instruct model. This is consistent with dataset designs: Fakepedia variants are evidence-centric and often adversarial/conflicting, pushing the model to prefer the provided context; StrategyQA/OpenBookQA rely more on commonsense priors and sparse cues, which encourages parametric recall. To understand the causal reason behind this knowledge interaction behaviour for these datasets, we investigate distribution of different knowledge interactions in Fig. 6. We find that both BaseFakepedia and MultihopFakepedia contain more conflicting examples than other knowledge interaction types (as defined in §3.2). Prior works (Cheng et al., 2024; Tao et al., 2024) suggest that for ‘conflicting’ examples, models tend to suppress PK when sufficient and relevant information is present in CK and for ‘supporting’ examples, models rely more on PK, with CK acting as a regularizer. Also, Tao et al. (2024) suggests that parametric recall is the default unless explicitly overridden by context. Overall, we conclude that for conflicting examples, the model aligns more with the CK direction and for supportive examples, the model aligns more with the PK direction in the rank-2 projection subspace, supporting observations from prior works.
@@
To understand the knowledge interaction dynamics during NLE generation (prompt template is described in Tab. 12 in §A.1.2), we analyze variation in ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ over all sequence steps of NLE generation for different knowledge interaction scenarios. Fig. 7 shows that for all datasets, during most of the NLE generations, the model starts with a higher CK, then considers both PK and CK with slight prioritisation of PK. However, for longer NLEs, CK and PK compete with each other with higher fluctuation. Longer NLEs indicate difficult examples with higher depth in multi-hop reasoning and higher token uncertainty (from Fig. 8), which force the model to iteratively reconcile PK with CK, resulting in this fluctuating behaviour.
@@
Figure 10: Comparison in individual PK-CK contribution in generating the answer token a for all the datasets from Meta-Llama3.1-8B-Instruct model between CoT and standard prompting.
@@ 4.4 RQ3: Can We Find Reasons for Hallucinations Based on PK-CK Interactions?
To characterize the knowledge interaction dynamics during the NLE generation in terms of context faithfulness, we investigate the knowledge alignments of hallucinated vs non-hallucinated responses in the rank-2 projection subspace. We utilize two RAG hallucination datasets: RAGTruth and Dollo (AC) Sun et al. (2025), of sizes $18240$ and $297$ examples respectively, containing examples from QA, Summarisation and Information Extraction. Each dataset provides human-annotated spans indicating hallucinated content across responses from multiple models. Importantly, whether a span is labeled as hallucinated is model-dependent: the same RAG input may yield hallucinated text for one model but not for another. Due to the limitations in the number of models covered in the two datasets, we only consider the data split corresponding to the Meta-Llama-3.1-8B-Instruct model. Fig. 9 illustrates the PK-CK knowledge interaction dynamics. The gap between PK and CK is much higher for the examples with hallucinated spans than for the examples with no hallucinated spans across the sequence steps. This result also aligns with similar observations of positive correlation of PK and hallucination in Sun et al. (2025).
@@ 5 Discussion
Future Directions and Broader Implications. These observations open several future directions. Extending subspace-based probing to other generative tasks such as summarization, dialogue, and retrieval-augmented reasoning, could reveal whether similar interaction dynamics generalize beyond NLEs. Integrating controllable subspace steering into model training or inference may enable fine-grained modulation of PK–CK balance, enhancing both interpretability and factual consistency. Finally, combining this approach with causal interventions such as neuron-level patching could illuminate how specific layers or modules mediate knowledge integration inside LLMs.
@@ 6 Conclusion
This work establishes that Parametric Knowledge (PK) and Context Knowledge (CK) interaction in LLMs is fundamentally multidimensional, not a binary choice between knowledge sources. Our rank-2 projection framework reveals that explanation (NLE) generation involves dynamic coordination between PK and CK, with their geometric balance serving as a direct indicator of factual reliability – hallucinated NLEs align strongly with the PK axis, while faithful NLEs balance both sources. This provides a mechanistic, internal-based signal for detecting hallucination in generated sequences. Beyond NLE generation, our framework can be used in the future to illuminate how models balance between internal knowledge and external grounding across diverse generative tasks. By enabling fine-grained multi-step tracking of knowledge contributions, this approach enables controllable steering of PK-CK dynamics, with implications for improving both interpretability and factual consistency in LLMs.
@@ A.1 Replication Details
where $\mathbf{H}_{w_{p}}\in\mathbb{R}^{n\times d}$, and $\mathbf{H}_{w_{c}}\in\mathbb{R}^{n\times d}$ are matrices of hidden representation of answer tokens $a(q,\varepsilon)$ and $a(q,c)$ guided by intents $w_{p}$ and $w_{c}$ respectively over $n$ examples.
@@ A.1.2 Prompt Template
Prompt template to generate controlled answers $a(q,\varepsilon)$, $a(q,c)$ and $a$ driven by intents $w_{p}$, $w_{c}$, and $w_{b}$ are described in Tab. 1. Prompt template for generating NLEs and CoT-based prompting over the datasets and models is described in Tab. 1.
@@
Answer the following query considering the provided context. Generate your final answer with only one word. If you are unable to answer the query, generate your final answer as "Unknown". Also, generate an explanation to determine your final answer. Return your output in JSON format: {"explanation": "your explanation here", "answer": "your final response here"}. Only include the JSON object in your response.
@@ C. Unknown
Answer the following query considering the provided context. Generate your final answer with only one word. If you are unable to answer the query, generate your final answer as "Unknown". Also, generate an explanation to determine your final answer. Return your output in JSON format: {"explanation": "your explanation here", "answer": "your final response here"}. Only include the JSON object in your response.
@@ A.1.3 Hyperparameters
We identify important layers $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ to capture individual PK and CK contributions, respectivel,y from the final answer $a$ via Patchscope using the hyperparameters described in Tab. 12.
@@ A.2 Additional Results
Figure 13: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four question–answer datasets using the Llama-3.1-8B-Instruct model. The split noise denotes cases where answers from individual knowledge sources agree with each other but differ from the final answer, i.e., $a(q,c)=a(a,\varepsilon)$ and $a(a,c)\neq a$.
Figure 14: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four question–answer datasets using the gemma-2-9b-it model. The split none denotes cases where answers from individual knowledge sources agree with each other but differ from the final answer, i.e., $a(q,c)=a(a,\varepsilon)$ and $a(a,c)\neq a$.
@@
Figure 15: Patchscope on OpenBookQA dataset from gemma-2-9b-it. a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$ results in a higher contribution of PK in generating the final answer, as the probability gap between the source and target is higher. b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$ results in a lower contribution of CK in generating the final answer, as the probability gap between the source and target is lower. We consider the common layers from both activation patching to learn the rank-2 projection subspace.
@@
Figure 16: Patchscope on OpenBookQA dataset from Mistral-7B-Instruct-v0.3. a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$ results in a higher contribution of PK in generating the final answer, as the probability gap between the source and target is higher. b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$ results in a lower contribution of CK in generating the final answer, as the probability gap between the source and target is lower. We consider the common layers from both activation patching to learn the rank-2 projection subspace.

### v2

@@ 4.1 Experimental Setting
Dataset and Model: We conduct experiments on three open-weight decoder-only instruct-based LMs: Llama-3.1-8B (Meta-Team, 2024), Gemma-2 9B (Gemma-Team, 2024) and Mistral-v0.3 7B (Jiang et al., 2023) using four publicly available QA datasets: BaseFakepedia, MultihopFakepedia (Minder et al., 2025), StrategyQA (Geva et al., 2021), and OpenBookQA (Mihaylov et al., 2018; Cheng et al., 2024). BaseFakepedia is a knowledge conflict dataset containing queries with 23 relations from Wikipedia. MultihopFakepedia is an extension of BaseFakepedia that contains queries that require extra-hop reasoning to generate answers. Both BaseFakepedia and MultihopFakepedia datasets are utilized for the rank-1 subpace analysis in Minder et al. (2025). StrategyQA is a multi-hop reasoning QA dataset that includes implicit reasoning steps in the queries. OpenBookQA is a commonsense reasoning-based QA dataset. Both StrategyQA and OpenBookQA datasets include ‘supportive’, ‘complementary’, ‘conflicting’, ‘irrelevant’, and ‘knowledge suppression’ PK-CK interactions, illustrated in Cheng et al. (2024).
@@
Explained Variance (Cumulative): Consider the matrix $H=[\{{\vec{\mathbf{h_{a_{j}}}}}\}_{j=1}^{N}],H\in\mathbb{R}^{N*d}$ as the concatenation of hidden representation $\vec{\mathbf{h_{a}}}\in\mathbb{R}^{d}$ of the answer token $a$ over $N$ examples. The $d$ singular values $\sigma_{1}>>\sigma_{2},...>>\sigma_{d}$ from the diagonal matrix $\Sigma\in\mathbb{R}^{d*d}$ after the Singular Value Decomposition (SVD) of $HH^{T}=A\Sigma B$ indicates the strength of $HH^{T}$ in $d$ orthonormal directions. For top-$r$ singular values $\sigma_{1}>>\sigma_{2},...>>\sigma_{r}$, the cumulative explained variance $EV_{r}$: $EV_{r}=\frac{\sum_{j=1}^{r}\sigma_{j}^{2}}{\sum_{j=1}^{d}\sigma_{j}^{2}}$
@@
Figure 2 shows, for each dataset, the distribution of the rank-1 subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h_{a}}}\rangle$, where $\vec{\mathbf{h_{a}}}$ is the hidden state of the answer token and $\vec{\mathbf{u}}$ is the learned rank-1 interaction direction. Each colored curve corresponds to one knowledge interaction type and shows how strongly answer representations align with $\vec{\mathbf{u}}$ (x-axis); the y-axis denotes the relative frequency, estimated via kernel density smoothing.
Across all datasets and knowledge interaction types (§3.2), the mean of the subspace component distribution converges to 0, indicating significant contribution from the orthogonal complement $\mathbf{(I-P)}$, showing insufficiency of the rank-1 projection subspace $\mathbf{P}$ in encoding the PK-CK interaction. This observed behaviour for the ‘complementary’ and ‘supportive’ interactions supports our hypothesis that, for examples where both PK and CK contribute equally, the rank-1 $\mathbf{P}$ fails to distinctly encode individual contributions. However, surprisingly, we observe a similar behaviour for the ‘conflicting’ interaction type as well, illustrating that $\mathbf{P}$ fails to also differentiate where the conflicts arise from. We observe similar results for the other two models (see Section A.2). These results indicate that different knowledge interactions are poorly captured by the rank-1 projection subspace, with most of the interaction signal residing in the orthogonal complement $\mathbf{(I-P)}$, thereby suggesting the necessity of a higher-rank subspace to effectively disentangle PK–CK contributions.
To understand the minimum rank required for the learnt projection subspace where the orthogonal complement contribution is minimum and it encodes different knowledge interaction types, we plot the cumulative explained variance $EV_{r}$ for different ranks across all datasets and models in Figure 3. $EV_{r}$ reaches $1.0$ at rank-2 and converges thereafter for all datasets and models.
In this section, we investigate the dynamics of individual PK and CK contribution ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ at every sequence step $i$ over the NLE $\mathcal{E}$ generation. Among the four datasets, OpenBookQA shows the largest jump in cumulative explained variance from rank-1 to rank-2 across all models (Figure 3). We therefore use OpenBookQA as our development set to (i) locate intent-sensitive layers via activation patching and (ii) learn the rank-2 projection subspace. We then freeze these choices and evaluate the resulting rank-2 subspace on all datasets. We also utilize the StrategyQA dataset and found that the learned parametric and contextual directions are highly stable, with average directional similarity exceeding $0.9$ for both axes (§A.3). We begin by identifying important layers for encoding individual knowledge contributions to generate the final answer in the rank-2 projection subspace. Figure 4 shows that patching from $w_{b}\rightarrow w_{p}$ (BOTH→PRI) (L13–18) yields a larger probability gain than $w_{b}\rightarrow w_{c}$ (BOTH→CTX) (L15–17), indicating that Llama-3.1-8B-Instruct relies more on PK than on CK for OpenBookQA. We observe similar results for gemma-2-9b-it and Mistral-7B-Instruct-v0.3 (see Figures 13 and 14). This PK dominance in a context-rich task suggests that the model tends to recall commonsense information from memory rather than grounding answers in the provided context, highlighting potential for future work on context-sensitive knowledge control.
Figure 2: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four QA datasets using Mistral-7B-Instruct-v0.3.
@@
Figure 4: Patchscope on OpenBookQA from Meta-Llama-3.1-8B-Instruct. (a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$. (b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$.
(a) Individual PK–CK contribution in generating the answer token.
(b) Distribution of knowledge interaction types across datasets.
Figure 5: Answer-level PK–CK alignment and distribution of knowledge interaction types across datasets.
Figure 6: PK-CK interaction dynamics over the sequence steps for Meta-Llama-3.1-8B-Instruct. The dotted red and blue lines indicate the mean and mode of NLE lengths. Figure 15 contains PK-CK dynamics for different interaction scenarios.
@@
To identify the individual knowledge contribution in generating the final answer token $a$, we project its hidden representation $\vec{\mathbf{h_{a}}}$ to the rank-2 projection subspace to compute the ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ as the PK-CK contributions. Figure 5(a) indicates an overall higher CK contribution for the BaseFakepedia and MultihopFakepedia, and a higher PK contribution for the StrategyQA and OpenBookQA for the Meta-Llama-3.1-8B-Instruct model. This is consistent with dataset designs: Fakepedia variants are evidence-centric and often adversarial/conflicting, pushing the model to prefer the provided context; StrategyQA/OpenBookQA rely more on commonsense priors and sparse cues, which encourage parametric recall. To understand the reason behind this knowledge interaction behaviour for these datasets, we investigate the distribution of different knowledge interactions in Fig. 5(b). We find that both BaseFakepedia and MultihopFakepedia contain more conflicting examples than other knowledge interaction types (as defined in §3.2). Prior works (Cheng et al., 2024; Tao et al., 2024) suggest that for ‘conflicting’ examples, models tend to suppress PK when sufficient and relevant information is present in CK and for ‘supporting’ examples, models rely more on PK, with CK acting as a regularizer. Also, Tao et al. (2024) suggests that parametric recall is the default unless explicitly overridden by context. Overall, we conclude that for conflicting examples, the model aligns more with the CK direction and for supportive examples, the model aligns more with the PK direction in the rank-2 projection subspace, supporting observations from prior works.
@@
To understand the knowledge interaction dynamics during NLE generation (prompt template is described in Tab. 10 in §A.1.2), we analyze variation in ${\alpha_{i}}^{p}$ and ${\alpha_{i}}^{c}$ over all sequence steps of NLE generation for different knowledge interaction scenarios. Figure 6 shows that for all datasets, during most of the NLE generations, the model starts with a higher CK, then considers both PK and CK with slight prioritization of PK. However, for longer NLEs, CK and PK compete with each other with higher fluctuation. Longer NLEs indicate difficult examples with higher depth in multi-hop reasoning and higher token uncertainty (from Figure 16), which force the model to iteratively reconcile PK with CK, resulting in this fluctuating behavior.
@@
Figure 8: Comparison in individual PK-CK contribution in generating the answer token $a$ for all the datasets from Meta-Llama-3.1-8B-Instruct model between CoT and standard prompting.
@@ 4.4 RQ3: Can We Find Reasons for Hallucinations Based on PK-CK Interactions?
To characterize the knowledge interaction dynamics during the NLE generation in terms of context faithfulness, we investigate the knowledge alignments of hallucinated vs non-hallucinated responses in the rank-2 projection subspace. We utilize two RAG hallucination datasets: RAGTruth and Dollo (AC) (Sun et al., 2025), of sizes $18240$ and $297$ examples respectively, containing examples from QA, Summarisation and Information Extraction. Each dataset provides human-annotated spans indicating hallucinated content across responses from multiple models. Importantly, whether a span is labelled as hallucinated is model-dependent: the same RAG input may yield hallucinated text for one model but not for another. Due to the limited number of models covered in the two datasets, we consider only the data split corresponding to the Meta-Llama-3.1-8B-Instruct model. Fig. 7 illustrates the PK-CK knowledge interaction dynamics. The gap between PK and CK is much higher for the examples with hallucinated spans than for the examples with no hallucinated spans across the sequence steps. This result also aligns with similar observations of positive correlation of PK and hallucination in Sun et al. (2025).
@@ 6 Conclusion
This work shows that the interaction between Parametric Knowledge (PK) and Context Knowledge (CK) in LLMs is inherently multidimensional rather than a binary choice. We introduce a rank-2 projection framework that (i) resolves the non-identifiability of rank-1 probes, (ii) disentangles token-level PK and CK components, and (iii) enables the first systematic multi-step analysis of PK–CK dynamics during Natural Language Explanation (NLE) generation. Across four QA datasets and three open-weight instruction-tuned LMs, rank-1 subspaces fail to capture diverse interaction regimes, whereas rank-2 suffices to explain the observed variance. Our step-wise analysis further reveals consistent patterns: hallucinated generations align strongly with the PK axis, while context-faithful generations maintain a more balanced PK–CK alignment, and Chain-of-Thought prompting shifts generations toward CK by reducing PK reliance. Beyond NLE generation, this subspace-based analysis provides a general, model-internal signal for studying how LLMs balance internal recall and external grounding.
@@ A.1 Replication Details
where $\mathbf{H}_{w_{p}}\in\mathbb{R}^{n\times d}$, and $\mathbf{H}_{w_{c}}\in\mathbb{R}^{n\times d}$ are matrices of hidden representation of answer tokens $a(q,\phi)$ and $a(q,c)$ guided by intents $w_{p}$ and $w_{c}$ respectively over $n$ examples.
@@ A.1.2 Prompt Template
Prompt template to generate controlled answers $a(q,\phi)$, $a(q,c)$ and $a$ driven by intents $w_{p}$, $w_{c}$, and $w_{b}$ are described in Table 1. The prompt template for generating NLEs and CoT-based prompting over the datasets and models is described in Figure 10.
@@
Answer the following query considering the provided context. Generate your final answer with only one word. If you are unable to answer the query, generate your final answer as ”Unknown”. Also, generate an explanation to determine your final answer. Return your output in JSON format: {"explanation": "your explanation here", "answer": "your final response here"}. Only include the JSON object in your response.
@@ C. Unknown
Answer the following query considering the provided context. Generate your final answer with only one word. If you are unable to answer the query, generate your final answer as ”Unknown”. Also, generate an explanation to determine your final answer. Return your output in JSON format: {"explanation": "your explanation here", "answer": "your final response here"}. Only include the JSON object in your response.
@@
Give your answer by analyzing step by step.
@@ A.1.3 Hyperparameters
We identify important layers $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ to capture individual PK and CK contributions, respectively, from the final answer $a$ via Patchscope using the hyperparameters described in Figure 10.
@@ A.2 Additional Results
Figure 11: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four question–answer datasets using the Llama-3.1-8B-Instruct model.
Figure 12: Kernel Density Estimate (KDE) of the PK-CK subspace component $\langle\vec{\mathbf{u}}^{T},\vec{\mathbf{h}}_{i}\rangle$ across different knowledge interaction types for four question–answer datasets using the gemma-2-9b-it model.
@@
Figure 13: Patchscope on OpenBookQA dataset from gemma-2-9b-it. (a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$ results in a higher contribution of PK in generating the final answer, as the probability gap between the source and target is higher. (b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$ results in a lower contribution of CK in generating the final answer, as the probability gap between the source and target is lower. We consider the common layers from both activation patching to learn the rank-2 projection subspace.
@@
Figure 14: Patchscope on OpenBookQA dataset from Mistral-7B-Instruct-v0.3. (a) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow p)}$ results in a higher contribution of PK in generating the final answer, as the probability gap between the source and target is higher. (b) Activation patching on $\mathcal{D}_{w}^{(b\rightarrow c)}$ results in a lower contribution of CK in generating the final answer, as the probability gap between the source and target is lower. We consider the common layers from both activation patching to learn the rank-2 projection subspace.
(a) Overall
(b) Supportive
(c) Complementary
(d) Conflicting
A.3 Robustness Analysis
We evaluate the robustness of PatchScope localization and the learned rank-2 PK–CK projection subspace along four axes: layer sensitivity, prompt perturbations, context variation, and geometric stability.
PatchScope layer sensitivity. To ensure that PK–CK disentanglement is not tied to a single layer, we evaluate the rank-2 subspace across six PatchScope-identified layers of Meta-Llama-3.1-8B-Instruct (L13–L18). Patching accuracy is stable across $6$ layers L13-18, $[0.82,0.84,0.83,0.85,0.87,0.86]$, with a mean of $0.84\pm 0.018$ and peak performance at L17 ($0.87$). The resulting stability score of $0.98$ indicates that disentanglement consistently emerges across this mid-to-late layer block rather than being a layer-specific artifact.
Prompt perturbation robustness. We test invariance to surface-form variations via paraphrasing, synonym substitution, and word-order perturbations. End-to-end prediction agreement remains high (model-dependent), at $0.92$, $0.88$, and $0.85$ for paraphrase, synonym, and word-order changes, respectively. To assess geometric robustness, we independently learn projection subspaces from original and perturbed prompts and measure their cosine similarity, obtaining mean consistency scores of $0.95\pm 0.04$, $0.93\pm 0.05$, and $0.89\pm 0.08$. These results indicate that the learned subspace is largely insensitive to superficial prompt variations and primarily reflects semantic knowledge alignment.
Context variation sanity check. If the projection subspace captures genuine knowledge interaction, PK-CK contributions should shift with context availability. Varying context from minimal (no context) to medium (first 50 CK tokens) to maximal (full context) yields: $\texttt{minimal}:(\alpha^{p}{=}0.82,\alpha^{c}{=}0.24)$, $\texttt{medium}:(\alpha^{p}{=}0.65,\alpha^{c}{=}0.58)$, $\texttt{maximal}:(\alpha^{p}{=}0.85,\alpha^{c}{=}0.87)$. The increase in CK contribution from minimal to medium context reflects greater reliance on evidence as it becomes available, while maximal context strongly activates both knowledge sources, consistent with supportive and complementary interactions rather than pure suppression of PK.
Geometric stability and separation. The learned parametric and contextual directions are highly stable across resampling, with average directional similarity exceeding $0.9$ for both axes. Moreover, the two directions remain well separated: the mean cosine overlap between PK and CK axes is $0.00$ across bootstraps. This consistent orthogonality supports the identifiability motivation for a rank-2 formulation. Overall, these results demonstrate that PK-CK disentanglement is stable across layers, robust to prompt surface perturbations, responsive to context evidence, and geometrically well-conditioned.